Fermi National Laboratory
 
Operation Summary for April 2, 2012 - April 4, 2012
 

Monday April 2

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest experiments, and the FTBF T-992 experiment taking beam.

During the day shift, at 8:23 AM, Operations disabled beam to T-992. At 8:49 AM, the NuMI beam permit tripped off for the first of two times in about three hours due to a power supply’s (E:V118) output drifting low; everything reset. A Switchyard (SY) person began a SeaQuest intensity study at 9:40 AM. Operations established beam to T-992 at 10:41 AM. A Muon Ring person began a study at 10:44 AM. The MiniBooNE beam permit tripped off at 11:19 AM due to bad beam positions; it reset. Operations turned off all SY beam at 12:56 PM to allow a power supply person to work on a critical device. MI personnel began study at 1:04 PM. Operators prepared the Neutrino Area for access at 1:06 PM. The MI studies ended at 2:23 PM.

During the evening shift, the Muon Ring study ended at 4:30 PM. Operations reported a problem with a Meson power supply (F:MW1W); Operators investigated and replaced a control card. At 7:40 PM, the NuMI beam permit tripped off due to V118; everything reset.

Tuesday April 3

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-992.

During the midnight shift, at 2:41 AM, the NuMI beam permit tripped off due to V118; everything reset. Operations disabled the T-992 beam at 3:08 AM.

During the day shift, at 9:03 AM, Linac personnel began the switch from the H- Source to the I- Source. Operations completed a Booster injection tune at 10:07 AM. A Booster person began a multibatch study at 10:42 AM. Operations established beam to T-992 at 11:23 AM. The Neutrino Area NS2 LCW system alarmed on a high surge tank level at 11:49 AM; a Water Group person investigated. At 12:04 PM, the MiniBooNE beam permit tripped off for the first of two times in 37 minutes due to bad beam positions; it reset. Muon Ring personnel began a study at 12:26 PM. At 12:41 PM, two Linac RF stations (LRF1 and LRF5) tripped off along with many quadrupole magnet power supplies; Linac personnel responded, brought up everything but LRF1, and reported the trips were all due to an I- Source spark. Linac resumed sending beam at 12:46 PM. A MI person began a g-2 (g minus 2) study at 1:04 PM. The NuMI beam permit tripped off at 2:01 PM due to V118; everything reset.

During the evening shift, the Muon Ring study ended at 4:28 PM. The g-2 study ended at 4:54 PM. Operations disabled T-992 beam at 8:40 PM. T-992 and SeaQuest made accesses during this shift.

Wednesday April 4

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at midnight the I- Source tripped off; it recovered. The NuMI beam permit tripped off at 12:41 AM for the first of two times in six hours due to V118; everything reset. At 3:17 AM, a Linac Klystron RF station (KRF4) wandered out of phase due to a cooling problem; Operators had the station back online by 3:39 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/04/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory